OpenStudy (anonymous):

0.2-1.5b+6.9-3a a=-5 and b=6

OpenStudy (anonymous):

replace a with -5 replace b with 6 0.2 -1.5(6) +6.9 -3(-5) multiply contents of parenthesis with coefficients, combine terms

OpenStudy (anonymous):

Ok I already substituted the variables for 6 and 3 so now I multiply the 1.5 by 6 and -5?

OpenStudy (anonymous):

Replace A for -5 and b to 6= 0.2 - (1.5*6) + 6.9 - (3*-5) 0.2 - (9 ) + 6.9 - (-15) = 13.1

OpenStudy (anonymous):

So is 31.1 our answer?

OpenStudy (anonymous):

Sorry I don't understand how you got 31.1

OpenStudy (anonymous):

do you need further explanation

OpenStudy (anonymous):

Yes please thank you

OpenStudy (anonymous):

Ok open your calculator or google calculator

OpenStudy (anonymous):

put in 0.2

OpenStudy (anonymous):

then the minus sign use open parentheses then - 9 close parenthesis

OpenStudy (anonymous):

ok do you mean to put this in the calculator? Because I already have this on paper

OpenStudy (anonymous):

what ever you want

OpenStudy (anonymous):

1.5 multiplied by 6 gives you 9

OpenStudy (anonymous):

-3 multiplied by a which is -5 = 15

OpenStudy (anonymous):

oh wow i just realized my mistake sorry Replace A for -5 and b to 6= 0.2 - (1.5*6) + 6.9 - (3*-5) 0.2 - (9 ) + 6.9 - (15) = -16.9

OpenStudy (anonymous):

Sorry my computer is slow something is wrong with it

OpenStudy (anonymous):

Its ok

OpenStudy (anonymous):

btw i got the correct answer i just accidentally put -15 before instead of 15 sorry

OpenStudy (anonymous):

I am sorry how do you get 16.9 I tried to put the equation in my calculator and got a completely different answer please explain

OpenStudy (anonymous):

ok so if you put 0.2 in then - 9 then add 6.9 - then you put - 15

OpenStudy (anonymous):

you should get -16.9

OpenStudy (anonymous):

tell me what you get

OpenStudy (anonymous):

as an answer

OpenStudy (anonymous):

oh ok I got it !!! now what?

OpenStudy (anonymous):

what did you get

OpenStudy (anonymous):

oh so this is our answer right?

OpenStudy (anonymous):

-16.9 is the answer

OpenStudy (anonymous):

Hope i helped

OpenStudy (anonymous):

I got-16.9

OpenStudy (anonymous):

good thats the answer

OpenStudy (anonymous):

you did thank You So much!! :)

OpenStudy (anonymous):

No problem any time